# Vectorized Computation of Euler Characteristic Functions and Transforms

### v1

@@ 5 Discussion
In this work, we presented a vectorized framework for computing WECFs (including ECFs and the WECT). This framework allows for efficient computation of these functions using tensor operations, which are well-suited for modern GPU hardware. Furthermore, our method works in full generality for weighted simplicial complexes (or cubical complexes) of arbitrary dimension with no structural constraints. Our framework is implemented in a publicly available Python package called pyECT. Experimentally, pyECT demonstrates a significant speedup over existing methods across a variety of image datasets. The implementation is also incredibly concise as the core vectorized operations leverage PyTorch, a well-maintained and optimized library for tensor computation. The core WECT algorithm in the pyECT implementation is only about $25$ lines of code, which although we acknowledge it is not a perfect metric, it is a testament to the simplicity of the approach.
@@
Another benefit of the approach presented in this work is that vectorized operations support automatic gradient computations. Potential future work includes exploring the use of this framework in topological machine learning applications and, in particular, incorporating topological transforms into neural network architectures. Additionally, we believe that this work will encourage further investigation into vectorized methods for computing other topological functions and transforms.
Acknowledgments
Computational efforts were performed on the Tempest High Performance Computing System, operated and supported by University Information Technology Research Cyberinfrastructure (RRID:SCR_026229) at Montana State University. ChatGPT was used in a brain-storming and debugging capacity for this work.

### v2

@@ 5 Discussion
In this work, we present a vectorized framework for computing WECFs (including ECFs and the WECT). This framework allows for efficient computation of these functions using tensor operations, which are well-suited for modern GPU hardware. Furthermore, our method works in full generality for weighted simplicial complexes (or cubical complexes) of arbitrary dimension with no structural constraints. Our framework is implemented in a publicly available Python package called pyECT. Experimentally, pyECT demonstrates a significant speedup over existing methods across a variety of image datasets. The implementation is also incredibly concise as the core vectorized operations leverage PyTorch, a well-maintained and optimized library for tensor computation. The core WECT algorithm in the pyECT implementation is only about $25$ lines of code, which we acknowledge is not a perfect metric, but is a testament to the simplicity of the approach.
@@
Another benefit of the approach presented in this work is that vectorized operations support automatic gradient computations. Potential future work includes exploring the use of this framework in differentiable forms of topological transforms (as in [17]) and, in particular, incorporating these transforms into neural network architectures. Additionally, we believe that this work will encourage further investigation into vectorized methods for computing other topological functions and transforms.
Acknowledgements
We thank Jack Ruder and Jacob Sriraman for their valuable assistance in developing and refining portions of the methodology and code used in this work. Computational efforts were performed on the Tempest High Performance Computing System, operated and supported by University Information Technology Research Cyberinfrastructure (RRID:SCR_026229) at Montana State University.
@@ Appendix A Examples of Tensor Operations
Appendix B Complexity Analysis
Here, we analyze the asymptotic complexity of Algorithm 1. Assume that the ambient dimension $n$ is fixed. Let $k$ be the total number of simplices of $K$, let $m$ be the number of vertex filters, and let NumVals be the number of values to sample the vertex filters over.
Next, the operation $\texttt{ScatterAdd}(\texttt{VIndices}^{T},\texttt{VertexWeights})$ is applied to diffWECFs. ScatterAdd adds a value from VertexWeights to diffWECFs for each entry of $\texttt{VIndices}^{T}$, resulting in a time complexity of $\Theta(km)$ and no additional memory usage as the operation is performed in place.
Now, for a fixed dimension $i\in\{1,2,\ldots n\}$, the tensor
$\texttt{SimpIndices}=\texttt{VIndices}[\texttt{SimplexVertices}]$
which takes $\Theta(km)$ time and memory.
Thus, we have a total time and memory complexity of $\Theta(nkm)=\Theta(km)$ after conclusion of the for-loop. Finally, a cumulative sum is applied to diffWECFs, taking $\Theta(m\texttt{NumVals})$ time and using no additional memory. This results in $\Theta(m(k+\texttt{NumVals}))$ for Algorithm 1 as the final time and memory complexity. In practice, $m$ and NumVals are typically held constant. Indeed, this is necessary to achieve consistent vectorizations of data. Holding $m$ and NumVals constant results in a theoretically optimal final time and memory complexity of $\Theta(k)$ for Algorithm 1.